OpenStudy (anonymous):

Newton Method HELP!!! numerical analysis establish newton's iterative scheme in simplified form,not involving the reciprocal of x, for the function f(x)=xR-x^-1

OpenStudy (anonymous):

what does \(xR-x^{-1}\) mean?

OpenStudy (anonymous):

R is the root

OpenStudy (anonymous):

root of?

OpenStudy (anonymous):

\[\sqrt{x}-\frac{1}{x}\]?

OpenStudy (anonymous):

in any case you want \[x_{n+1}=x_{n}-\frac{f(x_n)}{f'(x_n)}\] and then probably some algebra to make it look nicer